OpenStudy (anonymous):

help with connections algebra 1 A 9th plz

OpenStudy (anonymous):

Its algebra 1A and I need To Be Finished By 4 Today,Plzzz help

sam (.sam.):

Do you have a question?

sam (.sam.):

on maths*

OpenStudy (anonymous):

yes,but its quite a few of them,can you help

sam (.sam.):

Just post them

OpenStudy (anonymous):

1. What is an expression for the sale price of a bracelet that has been discounted 60% of its sticker price? Evaluate the expression for sticker prices of $90. Use the variable s for the price and p for the sticker price. 1.) s = p - 0.6p; $36 2.) s = p - 0.4p; $54 3.) s = p - 60p; 5,490 4.) s = p - 0.6p; $90

OpenStudy (anonymous):

well if you start with p and then want to subtract 60% off p, then you get (60% = 60/100 = 0.6) s = p - 0.6p.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so the correct answer is 1

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you can calculate it by plugging in p: 90 - 0.6*90 = 36

OpenStudy (anonymous):

ohh ok and I have a few more can you help??